OpenStudy (dengeki_daisy):

HELP PLEASE!! Which is greater, the total volume of three spheres, each of which has a diameter 3 in., the volume of one sphere that has a diameter 8 in., or the volume of a hemisphere with a diameter of 9 in.? a) The volume of three spheres, each with a diameter of 3 inches, is greater. b) The volume of one sphere with a diameter of 8 inches is greater. c) The volume of the hemisphere with a diameter of 9 in. is greater. d) The two volumes are the same.

OpenStudy (divyankasc):

The answer is c

OpenStudy (dengeki_daisy):

really?

Directrix (directrix):

1) Volume of Sphere is 4/3 * pi * r^3 If the diameter is 3, then the radius is 3/2 V = 4/3 * pi * (3/2)^3 V = 9*pi/2 Three of these spheres would have a volume of 3 * (9*pi/2) = 27*pi / 2

Directrix (directrix):

the volume of one sphere that has a diameter 8 in. V = 4/3 * pi * 4^3 where the radius is 4. V = 256* pi /3

OpenStudy (dengeki_daisy):

can you help with another?

Directrix (directrix):

the volume of a hemisphere with a diameter of 9 in. Volume of hemisphere = 1/2 that of the sphere V = 1/2 * 4/3 * pi * (9/2)^3 V = 343*pi/4

Directrix (directrix):

@dengeki_daisy Of these three volumes, which option is correct?

Directrix (directrix):

>The answer is c @Divyankasc I don't agree. Please show your work. Reminder: Direct Answers are against the OpenStudy rules.

OpenStudy (dengeki_daisy):

@Directrix I'm not really sure how to figure it out, can you explain it differently

Directrix (directrix):

No, I explained it the only way I know how. Learn the formula for the volume of a sphere and use it to compare volumes.

OpenStudy (dengeki_daisy):

okay I will

Directrix (directrix):

It comes down to which of these is the largest: 27*pi/2 or 256*pi/3 or 343* pi/4

OpenStudy (dengeki_daisy):

ohh well then, C

OpenStudy (dengeki_daisy):

@Directrix

Directrix (directrix):

Option C is the hemisphere and has volume V = 343*pi/4 .That approximates to 269.39. Is that larger than the approximation for the other two volumes?